how mole fraction is to use distinguished between solute and solvent in a solution when contain equal amount of solute and solvent
1
Expert's answer
2020-08-09T08:42:11-0400

 solute and solvent in a solution when contain equal amount of solute and solvent .


then Both Have Equal Mass .


let mass of solute and solvent both have " W " .


Molar Mass Of Solute And Solvent = M" AND M"" Respectevely .


 Moles of solute = W / M".


Moles of sovent = W / M"".


Moles of solution = W/ M" + W / M"".



SO ,MOLE FRACTION OF SOLUTE = MOLES OF SOLUTE / MOLESOF SOLUTION .


= W / M" / W/ M" + W / M"".


SO , MOLE FRACTION OF SOLVENT = MOLES OF SOLVENT / MOLESOF SOLUTION .

= W / M"" / W/ M" + W / M"".